# Unsupervised Density Neural Representation for CT Metal Artifact Reduction

### v1

@@ III-B1 MLP-based Parameterization
Without loss of generality, we present here our Diner model based on a 2D CT acquisition system for simplicity. However, the proposed approach can be easily extended to more advanced CT protocols (e.g., 3D cone-beam and helical CT) through a flexible geometry implementation. Given a 2D CT acquisition system, we represent the density map to be reconstructed as a function of spatial coordinates, defined by
@@ III-B2 Differentiable Forward Model
  • Performing our LACs decomposition model (Eq. (9)) to transform the MLP-predicted, energy-independent densities $\sigma(\mathbf{x})=g_{\phi}(\mathbf{x})$ into the polychromatic LACs $\mu(\mathbf{x},E)$ at multiple energy levels;
  • Using the discrete version of the polychromatic model (Eq. (3)) to generate the estimated measurement $\hat{p}(\mathbf{r})$ from the polychromatic LACs $\mu(\mathbf{x},E)$.
@@
$\hat{p}(\mathbf{r})=-\ln\sum_{i=1}^{N}\eta(E_{i})\cdot\text{exp}\left(-\sum_{\mathbf{x}\in\mathbf{r}}\mu(\mathbf{x},E_{i})\cdot\Delta\mathbf{x}\right),$ (11)
@@ III-B3 Loss Function
By leveraging the forward model $\mathcal{T}$, the MLP-predicted densities $\sigma(\mathbf{x})=g_{\phi}(\mathbf{x}),\forall\mathbf{x}\in\mathbf{r}$ are transformed into the CT measurement data $\hat{p}(\mathbf{r})$. Then, we define a loss function $\mathcal{L}_{\text{DC}}$ to compute the mean absolute error between the estimated $\hat{p}$ and real measurements ${p}$. The loss function can provide data consistency with the real measurements $p$. It can be expressed as below:
$\mathcal{L}_{\text{DC}}=\frac{1}{\left|\mathcal{R}\right|}\sum_{\mathbf{r}\in\mathcal{R}}\left|\hat{p}(\mathbf{r})-p(\mathbf{r})\right|,$ (13)
@@ III-C Virtual Monochromatic CT Image Reconstruction
After the optimization, we feed all spatial coordinates $\mathbf{x}$ into the well-trained MLP network to produce the corresponding densities $\sigma(\mathbf{x})=g_{\phi^{*}}(\mathbf{x})$, meaning a density map $\sigma(\cdot)$ can be reconstructed. Then, we use our LACs decomposition model (Eq. (9)) to generate the energy-dependent LAC maps $\mu(\cdot,E)$ at $N$ energy levels (i.e., polychromatic CT images). However, this work aims to recover a single monochromatic artifacts-free CT image. Consider a polychromatic X-ray composed of multiple monochromatic X-rays at $N$ energy levels, the energy level of its equivalent monochromatic X-ray is given by:
@@
Geometry Fan-beam Cone-beam Helical CT
@@
III-D Implementation Details
For our Diner, we employ the recent hash encoding [41] followed by two fully connected (FC) layers to implement the MLP network $g_{\phi}$. The hash encoding can effectively accelerate the training process. Its hyper-parameters are set as below: $L=16$, $T=2^{19}$, $F=8$, $N_{\mathrm{min}}=2$ and $b=2$. The two FC layers have 128 neurons, followed by ReLU and ELU activation functions, respectively. At each training iteration, we set the number of the random sampling X-rays as 80, i.e., $|\mathcal{R}|=80$ in Eq. (13). We use the Adam optimizer [45] with default hyperparameters. The learning rate starts from 1e-3 and decays by half per 500 epochs. The total training epochs is 2000, which takes about 1.2 minutes on a single NVIDIA RTX TITIAN GPU. Note that all hyper-parameters are tuned on 10 samples from the DeepLesion dataset [43] and then fixed across all other samples.
Figure 3: (a) The commercial Bruker SKYSCAN 1276 micro-CT scanner used in our mouse tight data acquisition and (b) The commercial UIH uCT 768 scanner used in our body phantom data acquisition.
@@ IV-A2 Mouse Tight Real-world Dataset
To evaluate the effectiveness of our Diner model in biomedical X-ray CT settings, we scan five mouse-thigh samples containing intramedullary needles of 304 stainless steel metal on a commercial Bruker SKYSCAN 1276 micro-CT scanner, as shown in Figure 3. The CT acquisition protocol is shown in Table I. We use the SPEKIT toolkit [42] to estimate the X-ray energy spectrum of the scanner. The Ethics Committee of ShanghaiTech University approves this data collection.
@@ IV-A3 Body Phantom Real-world Dataset
To test the MAR performance of the proposed method on clinical helical CT, we measure a human body phantom containing Ti metals using a commercial UIH X-ray CT scanner (uCT 768). To align with clinical settings, we use a standard clinical acquisition protocol, which is demonstrated in Table I. This data acquisition is approved by UIH Co., Ltd. and the Ethics Committee of ShanghaiTech University.
@@ IV-B Baselines and Evaluation Metrics
We compare the proposed Diner with eight representative MAR methods belonging to four categories: i) Three model-based methods (FBP [2], LI [4], and ART [6]); ii) Three supervised DL methods (CNN-MAR [7], DICDNet [10], and ACDNet [9]); iii) One unpaired-supervised DL method (ADN [11]); iv) One unsupervised DL method (Polyner [26]). It should be noted that the supervised and unpaired-supervised DL methods are evaluated based on the pre-trained models provided by the authors to ensure a fair comparison. Among them, CNN-MAR [7] is trained on the XCOM dataset [30], while DICDNet [10], ACDNet [9], and ADN [11] are trained on the DeepLesion dataset [43].
We use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) [46] as quantitative metrics and implement them by the Python library scikit-image [47].
@@ IV-C Comparison with Baselines on Simulation Data
Table II shows the quantitative results. Generally, our Diner produces the best performance for all cases. On the DeepLesion dataset [43], our unsupervised Diner not only significantly outperforms the unsupervised Polyner [26], but also slightly performs better than the two supervised methods (DICDNet [10] and ACDNet [9]). For example, the PSNR respectively improves by +1.94 dB, +1.24 dB, and +0.74 dB for the titanium metals. While on the LIDC dataset [44], the supervised ACDNet and ACDNet trained on the DeepLesion dataset [43] fail to achieve pleasant MAR performance due to the OOD problem. They even perform worse than the traditional model-based LI algorithm [4], such as -2.46 dB and -1.20 dB in the PSNR for the 304 stainless steels. In comparison, our Diner still yields the best MAR reconstructions, outperforming the second-best Polyner [26] by about +2.06 dB in PSNR for the 304 stainless steels. We show the qualitative results in Figure 4. Traditional FBP algorithm [2] and supervised CNN-MAR trained on the XCOM [30] cannot recover satisfactory results, including severe artifacts on both datasets. While ACDNet produces a clear but slightly smooth MAR reconstruction on the DeepLesion dataset [43], its result on the LIDC dataset suffers from severe shadow artifacts. In comparison, the two unsupervised methods (Polyner [26] and our Diner) achieve robust results across the two datasets. However, the results of Polyner [26] contain slight artifacts around the metal, while our Diner produces clean and fine-detailed MAR reconstruction.
@@ IV-D1 On Mouse Tight Dataset
We collect five mouse-thigh samples containing an intramedullary needle with a commercial Bruker SKYSCAN 1276 micro-CT scanner. The initial intent of the mouse model is to investigate the progress of bone-fracture healing in mice. The setup is shown in Figure 3. The intramedullary needle was implanted into the mouse tight one week before the CT scan. To further assess the accuracy of MAR CT reconstruction, the mouse thighs are scanned twice before and after needle removal. The reference image is the clear version of the mouse thigh sample after removing the intramedullary needle. We show the qualitative results in Figure 5. Compared with NRECON, a reconstruction toolkit developed by Bruker and equipped with Bruker SKYSCAN 1276 micro-CT scanner, our Diner produces a clear CT reconstruction closer to reference. Moreover, we use CTan, a bone analysis software developed by Bruker, to perform a bone analysis on these CT volumes. The results are shown in Figure 6. In terms of bone volume (BV), percent bone volume (BV/TV), and trabecular number (Tb.N), the reconstruction of NRECON obtain distinct higher scores than the clear reference image. This is because the metal artifacts in the CT volume are identified as high-density bone tissue by the CTan software. In comparison, our Diner effectively removes these metal artifacts and thus obtains very close scores to the reference image.
@@ IV-D2 On Body Phantom Dataset
We conduct scans on two versions (with and without a metal stick) of a body phantom using the UIH commercial uCT 768 scanner according to a clinical acquisition protocol (See Table I for detailed parameters). The experimental setup is depicted in Figure 3. The qualitative results are presented in Figure 7. Here, ”UIH uCT 768” denotes the reconstruction generated by the algorithm integrated into the clinical UIH uCT 768 scanner. Upon visual inspection of the CT slices from axial, sagittal, and coronal views (Rows 1–3), our Diner effectively reduces metal artifacts while finely preserving image details. Additionally, we utilize the Bee DICOM Viewer software for volume rendering of the CT reconstructions. These results are illustrated in Figure 7 (Rows 4–5). Also, the rendering outcome of our Diner is clear and fine-detailed, closely resembling that of the reference image. In summary, these ongoing experimental investigations conducted using the UHI uCT 768 scanner robustly demonstrate the efficacy of our Diner in clinical settings.
@@
TABLE III: Quantitative results of polychromatic CT images reconstructed by our Diner and Polynner [26] for three types of metals on the DeepLesion dataset [43].
Method Types of Metals PSNR SSIM
Polyner [26] Titanium 29.75$\pm$5.16 0.9349$\pm$0.0738
Chromium 29.71$\pm$5.14 0.9336$\pm$0.0736
Stainless Steel 29.54$\pm$5.00 0.9320$\pm$0.0732
Diner (Ours) Titanium 37.31$\pm$2.36 0.9820$\pm$0.0052
Chromium 37.05$\pm$2.28 0.9800$\pm$0.0056
Stainless Steel 36.92$\pm$2.24 0.9790$\pm$0.0059
@@ IV-E Reconstruction of Polychromatic Images
In our previous Polyner [26], the recovery of the polychromatic CT images is ill-posed on the spectrum distribution. In contrast, our Diner solves the energy-independent density map, which improves the reconstruction of the energy-dependent LACs. By conducting our LAC decomposition model (Eq. (9)), we can generate the LAC maps from the resolved density map. Here, we compare our Diner with Polyner for reconstructing the polychromatic images. Figure 8 show the solved polychromatic images at four energy levels $E=\{40,60,100,120\}$. From the visualization, we observe that the results of our Diner model are very close to the GT samples in terms of image structure and contrast. In comparison, the reconstructions of Polyner are satisfactory for image structure but inconsistent for image contrast, i.e., spectrum distribution. We show the average LAC curve of the polychromatic images over the X-ray energy in Figure 9. Polyner model produces an over-smooth spectrum distribution due to an energy-smooth regulation. In contrast, our Diner is highly consistent with GT, benefiting from the LAC decomposition model (Eq. (8)). Table III also shows the quantitative results of polychromatic images by Polyner [26] and Diner for all cases. The proposed Diner significantly improves the performance.
TABLE IV: Quantitative results of our Diner with different forward models on the DeepLesion dataset [43].
@@ IV-F1 Influence of Forward Model
Our Diner proposes a new forward model $\mathcal{T}$ (Eq. (12)) to enable the INR network to reconstruct the densities of the object. Here, we explore the influence of the forward model $\mathcal{T}$ on the MAR performance. Specifically, we replace it with the linear integral model (Eq. (3)) used in the existing SOTA MAR techniques [7, 4, 2, 11, 9, 10] and make all other model configurations the same for a fair comparison. Table IV shows the quantitative results. Our proposed nonlinear model $\mathcal{T}$ produces significant performance improvements compared with the conventional linear integral, where PSNR improves by +4.72 dB. We show the qualitative results in Figure 10. The linear model fails to tackle the shadow artifacts caused by the BHE. In comparison, the result of our nonlinear model is visually clear and very close to the GT sample. The experimental comparison indicates that our nonlinear model is critical for the MAR reconstructions.
TABLE V: Quantitative results of our Diner with different spectrum priors $\eta$ on the DeepLesion dataset [43].
@@ IV-F2 Influence of Energy Spectrum Prior
In our forward model $\mathcal{T}$, the X-ray energy spectrum $\eta\in\mathbb{R}^{N}$ is considered as available prior knowledge. Here, we study its influence on the model performance. Three types of spectra are compared. (1) Random Noise, a vector of length $N$ sampled from the Gaussian distribution $\mathcal{N}(0,1)$. (2) Noisy Spectrum, a spectrum corrupted by an additive noise following the Gaussian distribution $\mathcal{N}(0,\eta_{\mathrm{max}}\cdot c)$, where $\eta_{\mathrm{max}}$ denotes the maximum values of the clear spectrum, and $c$ is set to $3\%$ and $1.5\%$ to simulate two levels of noise. (3) Clear spectrum, i.e., a GT spectrum. Table V shows the quantitative results. There are two observations. First, using the random noise as the energy spectrum causes severe degradation of the MAR performance, which decreases by 16.74 dB in terms of PSNR compared to using the clear spectrum. Second, the use of the noisy spectra decreases the performance of our Diner, but at an acceptable level. For example, the PSNR reaches 35.64 dB and 36.80 dB for the noisy spectra (1.5% and 3%). Figure 11 shows the qualitative results. From the visualization, we observe that the result of the noisy spectrum is significantly better than that of the random noise. In summary, the energy spectrum is required for our Diner, but it does not significantly affect the MAR performance if inaccurate.

### v2

@@ III-B1 MLP-based Parameterization
Without loss of generality, we introduce Diner model based on 2D CT acquisition systems for simplicity here. However, the proposed technique can be easily extended to more advanced CT protocols (e.g., 3D cone-beam and helical CT) through a flexible geometry implementation. Given a 2D CT acquisition system, we represent the density map to be reconstructed as a function of spatial coordinates, defined by
@@ III-B2 Differentiable Forward Model
  • Performing our LAC decomposition model (Eq. 9) to transform the MLP-predicted, energy-independent densities $\sigma(\mathbf{x})=g_{\phi}(\mathbf{x})$ into the polychromatic LACs $\mu(\mathbf{x},E)$ at multiple energy levels;
  • Using the discrete version of the polychromatic model (Eq. 3) to generate the estimated measurement $\hat{\rho}(\mathbf{r})$ from the polychromatic LACs $\mu(\mathbf{x},E)$.
@@
$\hat{\rho}(\mathbf{r})=-\ln\sum_{i=1}^{N}\eta(E_{i})\cdot\text{exp}\left(-\sum_{\mathbf{x}\in\mathbf{r}}\mu(\mathbf{x},E_{i})\cdot\Delta\mathbf{x}\right),$ (11)
@@ III-B3 Loss Function
By leveraging the forward model $\mathcal{T}$, the MLP-predicted densities $\sigma(\mathbf{x})=g_{\phi}(\mathbf{x}),\forall\mathbf{x}\in\mathbf{r}$ are transformed into the CT measurement data $\hat{\rho}(\mathbf{r})$. Then, we define a loss function $\mathcal{L}_{\text{DC}}$ to compute the mean absolute error between the estimated $\hat{\rho}$ and real measurements ${\rho}$. The loss function can provide data consistency with the real measurements $\rho$. It can be expressed as below:
$\mathcal{L}_{\text{DC}}=\frac{1}{\left|\mathcal{R}\right|}\sum_{\mathbf{r}\in\mathcal{R}}\left|\hat{\rho}(\mathbf{r})-\rho(\mathbf{r})\right|,$ (13)
@@ III-C Virtual Monochromatic CT Image Reconstruction
After the optimization, we feed all spatial coordinates $\mathbf{x}$ into the well-trained MLP network to produce the corresponding densities $\sigma(\mathbf{x})=g_{\phi^{*}}(\mathbf{x})$, meaning a density map $\sigma(\cdot)$ can be reconstructed. Then, we use our LACs decomposition model (Eq. 9) to generate the energy-dependent LAC maps $\mu(\cdot,E)$ at $N$ energy levels (i.e., polychromatic CT images). However, this work aims to recover a single monochromatic artifacts-free CT image. Consider a polychromatic X-ray composed of multiple monochromatic X-rays at $N$ energy levels, the energy level of its equivalent monochromatic X-ray is given by
@@
III-D Implementation Details
For our Diner, we employ the recent hash encoding [41] followed by two fully connected (FC) layers to implement the MLP network $g_{\phi}$. The hash encoding can effectively accelerate the training process. Its hyper-parameters are set as below: $L=16$, $T=2^{19}$, $F=8$, $N_{\mathrm{min}}=2$ and $b=2$. The two FC layers have 128 neurons, followed by ReLU and ELU activation functions, respectively. At each training iteration, we set the number of the random sampling X-rays as 80, i.e., $|\mathcal{R}|=80$ in Eq. 13. We use the Adam optimizer with default hyperparameters. The learning rate starts from 1e-3 and decays by half per 500 epochs. The total training epochs is 2000, which takes about 1.2 minutes on a single NVIDIA RTX TITIAN GPU. Note that all hyper-parameters are tuned on 10 samples from the DeepLesion dataset [43] and then fixed across all other samples.
@@ III-D Implementation Details
Geometry 2D Fan-beam 3D Cone-beam 3D Helical
@@ IV-A1 DeepLesion and LIDC Simulation Datasets
Figure 3: a) A commercial Bruker SKYSCAN 1276 micro-CT scanner used in our mouse tight data acquisition and b) A commercial UIH uCT 768 scanner used in our body phantom data acquisition.
@@ IV-A2 Mouse Tight Real-world Dataset
To evaluate the effectiveness of our Diner model in biomedical X-ray CT settings, we scan five mouse-thigh samples containing intramedullary needles of 304 stainless steel metal on a commercial Bruker SKYSCAN 1276 micro-CT scanner, as shown in Figure 3a. The CT acquisition protocol is shown in Table I. We use the SPEKIT toolkit [42] to estimate the X-ray energy spectrum of the scanner.
@@ IV-A3 Body Phantom Real-world Dataset
To test the MAR performance of the proposed method on clinical helical CT, we measure a human body phantom containing Ti metals using a commercial UIH uCT 768 scanner (Figure 3b). To align with clinical settings, we use a standard clinical acquisition protocol, which is demonstrated in Table I.
@@ IV-B Baselines and Evaluation Metrics
We compare the proposed Diner with eight representative MAR methods belonging to four categories: i) Three model-based methods (FBP [2], LI [4], and ART [5]); ii) Three supervised DL methods (CNN-MAR [6], DICDNet [9], and ACDNet [8]); iii) One unpaired-supervised DL method (ADN [10]); iv) One unsupervised DL method (Polyner [26]). It should be noted that the supervised and unpaired-supervised DL methods are evaluated based on the pre-trained models provided by the authors to ensure a fair comparison. Among them, CNN-MAR [6] is trained on the XCOM dataset [34], while DICDNet [9], ACDNet [8], and ADN [10] are trained on the DeepLesion dataset [43]. We use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) as quantitative metrics.
@@ IV-C Comparison with Baselines on Simulation Data
Table II shows the quantitative results. Generally, our Diner produces the best performance for all cases. On the DeepLesion dataset [43], our unsupervised Diner not only significantly outperforms the unsupervised Polyner [26], but also slightly performs better than the two supervised methods (DICDNet [9] and ACDNet [8]). For example, the PSNR respectively improves by +1.94 dB, +1.24 dB, and +0.74 dB for the titanium metals. While on the LIDC dataset [44], the supervised ACDNet and ACDNet trained on the DeepLesion dataset [43] fail to achieve pleasant MAR performance due to the OOD problem. They even perform worse than the traditional model-based LI algorithm [4], such as -2.46 dB and -1.20 dB in the PSNR for the 304 stainless steels. In comparison, our Diner still yields the best MAR reconstructions, outperforming the second-best Polyner [26] by about +2.06 dB in PSNR for the 304 stainless steels.
We show the qualitative results in Figure 4. Traditional FBP algorithm [2] and supervised CNN-MAR trained on the XCOM [34] cannot recover satisfactory results, including severe artifacts on both datasets. While ACDNet produces a clear but slightly smooth MAR reconstruction on the DeepLesion dataset [43], its result on the LIDC dataset suffers from severe shadow artifacts. In comparison, the two unsupervised methods (Polyner [26] and our Diner) achieve robust results across the two datasets. However, the results of Polyner [26] contain slight artifacts around the metal, while our Diner produces clean and fine-detailed MAR reconstruction.
@@ IV-D1 On Mouse Tight Dataset
We collect five mouse-thigh samples containing an intramedullary needle with a commercial Bruker SKYSCAN 1276 micro-CT scanner. The initial intent of the mouse model is to investigate the progress of bone-fracture healing in mice. The setup is shown in Figure 3a. The intramedullary needle was implanted into the mouse tight one week before the CT scan. To further assess the accuracy of MAR CT reconstruction, the mouse thighs are scanned twice before and after needle removal. The reference image is the clear version of the mouse thigh sample after removing the intramedullary needle.
We show the qualitative results in Figure 5. Compared with NRECON, a reconstruction toolkit developed by Bruker and equipped with Bruker SKYSCAN 1276 micro-CT scanner, our Diner produces a clear CT reconstruction closer to reference. Moreover, we use CTan, a bone analysis software developed by Bruker, to perform a bone analysis on these CT volumes. The results are shown in Figure 6. In terms of bone volume (BV), percent bone volume (BV/TV), and trabecular number (Tb.N), the reconstruction of NRECON obtain distinct higher scores than the clear reference image. This is because the metal artifacts in the CT volume are identified as high-density bone tissue by the CTan software. In comparison, our Diner effectively removes these metal artifacts and thus obtains very close scores to the reference image.
@@ IV-D2 On Body Phantom Dataset
We conduct scans on two versions (with and without a metal stick) of a body phantom using the UIH commercial uCT 768 scanner according to a clinical acquisition protocol (See Table I for detailed parameters). The experimental setup is depicted in Figure 3b. The qualitative results are presented in Figure 7. Here, “UIH uCT 768” denotes the reconstruction generated by the algorithm integrated into the clinical UIH uCT 768 scanner. Upon visual inspection of the CT slices from axial, sagittal, and coronal views (Rows 1–3), our Diner effectively reduces metal artifacts while finely preserving image details.
Additionally, we utilize the Bee DICOM Viewer software for volume rendering of the CT reconstructions. These results are illustrated in Figure 7 (Rows 4–5). Also, the rendering outcome of our Diner is clear and fine-detailed, closely resembling that of the reference image. In summary, these ongoing experimental investigations conducted using the UHI uCT 768 scanner robustly demonstrate the efficacy of our Diner in clinical settings.
@@ IV-E Reconstruction of Polychromatic Images
In Polyner [26], the recovery of the polychromatic CT images is ill-posed on the spectrum distribution. In contrast, our Diner solves the energy-independent density map, which improves the reconstruction of the energy-dependent LACs. By conducting our LAC decomposition model (Eq. 9), we can generate the LAC maps from the resolved density map. Here, we compare our Diner with Polyner for reconstructing the polychromatic images. Figure 8 show the solved polychromatic images at four energy levels $E=\{40,60,100,120\}$. From the visualization, we observe that the results of our Diner model are very close to the GT samples in terms of image structure and contrast. In comparison, the reconstructions of Polyner are satisfactory for image structure but inconsistent for image contrast, i.e., spectrum distribution.
We show the average LAC curve of the polychromatic images over the X-ray energy in Figure 9. Polyner model produces an over-smooth spectrum distribution due to an energy-smooth regulation. In contrast, our Diner is highly consistent with GT, benefiting from the LAC decomposition model (Eq. 8). To sum up, our Diner greatly outperforms our previous Polyner model for the polychromatic CT images.
TABLE III: Quantitative results of our Diner with different forward models on the DeepLesion dataset [43].
@@ IV-F1 Influence of Forward Model
Our Diner proposes a new forward model $\mathcal{T}$ (Eq. 12) to enable the INR network to reconstruct the densities of the object. Here, we explore the influence of the forward model $\mathcal{T}$ on the MAR performance. Specifically, we replace it with the linear integral model (Eq. 3) used in the existing SOTA MAR techniques [6, 4, 2, 10, 8, 9] and make all other model configurations the same for a fair comparison. Table III shows the quantitative results. Our proposed nonlinear model $\mathcal{T}$ produces significant performance improvements compared with the conventional linear integral, where PSNR improves by +4.72 dB. We show the qualitative results in Figure 10. The linear model fails to tackle the shadow artifacts caused by the BHE. In comparison, the result of our nonlinear model is visually clear and close to the GT.
TABLE IV: Quantitative results of our Diner with different spectrum priors $\eta$ on the DeepLesion dataset [43].
@@ IV-F2 Influence of Energy Spectrum Prior
In our forward model $\mathcal{T}$, the X-ray energy spectrum $\eta\in\mathbb{R}^{N}$ is considered as available prior knowledge. Here, we study its influence on the model performance. Three types of spectra are compared. (1) Random Noise, a vector of length $N$ sampled from the Gaussian distribution $\mathcal{N}(0,1)$. (2) Noisy Spectrum, a spectrum corrupted by an additive noise following the Gaussian distribution $\mathcal{N}(0,\eta_{\mathrm{max}}\cdot c)$, where $\eta_{\mathrm{max}}$ denotes the maximum values of the clear spectrum, and $c$ is set to $3\%$ and $1.5\%$ to simulate two levels of noise. (3) Clear spectrum, i.e., a GT spectrum. Table IV shows the quantitative results. There are two observations. First, using the random noise as the energy spectrum causes severe degradation of the MAR performance, which decreases by 16.74 dB in terms of PSNR compared to using the clear spectrum. Second, the use of the noisy spectra decreases the performance of our Diner, but at an acceptable level. For example, the PSNR reaches 35.64 dB and 36.80 dB for the noisy spectra (1.5% and 3%). Figure 11 shows the qualitative results. From the visualization, we observe that the result of the noisy spectrum is significantly better than that of the random noise. In summary, the energy spectrum is required for our Diner, but it does not significantly affect the MAR performance if inaccurate.